Imagine a perfectly competitive firm producing good A with cost function

TC=400+20Q-2Q2+2/3Q3, where Q is quantity produced

a. determine the firm’s short run supply curve

b. What is the profit maximizing level of output when price of A is birr 180?

1
Expert's answer
2021-12-16T14:38:36-0500

a. the firm’s short run supply curve is MC

"MC=TC'=(400+20Q-2Q2+2\/3Q3)'=20-4Q+2Q^2"

b.

MC=P

"20-4Q+2Q^2=180"

"20-4Q+2Q^2-180=0"

"2Q^2-4Q-160=0"

"Q^2-2Q-80=0"

Q=10

"profit=MR-MC=180\times10-(20-4\times10+2(10^2))=1800-20+40-200=1620"